# Multi-Agent Probabilistic Ensembles with Trajectory Sampling for Connected Autonomous Vehicles

### v1

@@ III-B System Model
[…]
@@ IV-B Planning and Control with Learned Dynamics
Based on a learned dynamics model $\tilde{\mathcal{P}}^{(i)}$, MA-PETS tries to obtain a solution to (6) by resorting to a sample-efficiency controller MPC [23]. Generally speaking, MPC avoids cumbersome gradient computation and exhibits strong robustness to the aleatoric and epistemic uncertainty in the learned dynamics model. Without loss of generality, assume that for a time-step $t$, agent $i$ observes the state $s_{t}^{(i)}$. Afterwards, agent $i$ leverages the Cross-Entropy Method (CEM) [44] to generate $Q$ candidate action sequences within a $w$-horizon of MPC. Initially, each candidate action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)},\forall q\in[Q]$ is sampled following a Gaussian distribution $\mathcal{N}(\mu_{\phi},\sigma_{\phi})$ parameterized by $\phi$. Meanwhile, agent $i$ complements possible next-time-step states from the dynamics model ensembles, by adopting particle-based TS [45]. In other words, for each candidate action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)}$, TS predicts plausible state-action trajectories by simultaneously creating $P$ particles that propagate a set of Monte Carlo samples from the state $s_{t}^{(i)}$. Due to the randomness in the learned and assumed time-invariant dynamics model, each particle $p\in[P]$ can be propagated by $s_{t^{\prime}+1,p}^{(i)}={\tilde{\mathcal{P}}_{\theta_{b}}^{(i)}}\left(s_{t^{\prime},p}^{(i)},a_{t^{\prime}}^{(i)}\right)$ ($t^{\prime}\in[t,t+w-2]$), where $\tilde{\mathcal{P}}_{\theta_{b}}^{(i)},b\in[B]$ is a randomly selected dynamics model in the ensemble model. Therefore, the set of plausible state-action trajectories for each candidate action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)}=\left(a_{t}^{(q,i)},\ldots,a_{t+w-1}^{(q,i)}\right)$ ($q\in[Q]$) consists of $P$ parallel propagated states with the same action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)}$. Afterwards, given the calibrated known reward function $\mathcal{R}^{(i)}$, the evaluation of a candidate action sequence can be derived from the average cumulative reward of $P$ different parallel action-state trajectories, that is,
@@ V-A EVI for an Optimistic MDP
Notably, Assumption 1 holds naturally, if we neglect the possible approximation error of DNN in MA-PETS, as it has a trivial impact on understanding the sample efficiency incurred by information exchange. On the other hand, Assumption 2 could be easily met by intentionally ignoring the experienced visits of some “diligent” agents. Moreover, Assumption 3 implies unanimously scaling the group regret bound by $r_{\max}$, which does not affect learning the contributing impact from inter-agent communications.
@@
Consistent with UCRL2, we allow each agent $i$ to enter into a new episode $k+1$ once there exists a state-action pair $(s,a)$ that has just been played and satisfies $u_{k}^{(i)}(s,a)=N^{(i)+}_{k}(s,a)$. Here $u_{k}^{(i)}(s,a)$ denotes the number of state-action $(s,a)$ visits by agent $i$ in the episode $k$ and $N^{(i)+}_{k}(s,a)=\max\left\{1,N^{(i)}_{k}(s,a)\right\}$. By Assumption 2, all agents could enter into the next episode in a “simultaneous” manner, by intentionally ignoring the experienced visits of a “diligent” agent $i$ after meeting $u_{k}^{(i)}(s,a)=N^{(i)+}_{k}(s,a)$ in episode $k$. Therefore, at the very beginning of episode $k+1$, for each agent $i$, $N_{k+1}^{(i)}(s,a)=N_{k}^{(i)}(s,a)+u_{k}^{(i)}(s,a)$ for all state-action pairs$(s,a)$ This setting is similar to the doubling criterion in single-agent UCRL2 and ensures that each episode is long enough to allow sufficient learning. Furthermore, different from the single-agent UCRL2, at the end of an episode $k$, consistent with in MA-PETS, each agent $i$ has access to all the set of neighboring CAVs $\mathcal{F}_{d,k}^{(i)}$, so as to obtain the latest dataset $\mathcal{D}_{t_{k}-H:t_{k}}^{(j)}$, $\forall j\in\mathcal{F}_{d,k}^{(i)}$ and constitute $\mathcal{D}_{0:t_{k}}^{(i)+}$. Again, according to Assumption 2, $H=\sum_{s,a}u_{k}^{(i)}(s,a)$. Afterwards, the EVI proceeds to estimate the transition probability function $\mathcal{P}^{(i)}(\cdot\mid s,a)$. In particular, various high-probability confidence intervals of the true MDP $M^{(i)}$ can be conjectured according to different concentration inequalities. For example, based on Hoeffding’s inequality [46], the confidence interval for the estimated transition probabilities can be given as
@@
Correspondingly, a policy for an optimistic MDP can be attained from learning in the conjectured plausible MDPs, and then the impact of the communication range $d$ on the performance of MA-PETS can be analysed, in terms of the multi-agent group regret in (5).
@@ V-B1 Results for Single-Agent Regret Bound
Consider $M^{(i)}$ as a time-homogeneous and communicating MDP with a policy $\pi^{\prime\ast(i)}$. Meanwhile, by Lemma 1, the optimal $T$-step reward $\mathcal{R}^{\ast(i)}_{T}(M^{(i)})$ can be approximated by the optimistic average reward $\rho^{\ast(i)}$. Therefore, if $\pi^{\prime\ast}(i)$ is implemented in $M^{(i)}$ for $T$ steps, we have the following lemma.
@@
Under Lemma 1, the differences between the observed rewards $r_{t}^{(i)}$, acquired when agent $i$ follows the policy $\pi^{\prime\ast(i)}$ to choose action $a_{t}^{(i)}$ in state $s_{t}^{(i)}$ at step $t$, and the aforementioned optimistic average reward $\rho^{\ast(i)}$ forms a martingale difference sequence. Besides, on the basis of Hoeffding’s inequality, with a probability at least $1-\frac{\delta}{12(T)^{5/4}}$, the single-agent bound can be bounded as
@@
where $C_{1}=\frac{5}{8}$ is a constant, and we define the regret in a single episode $k$ as $\Delta^{(i)}_{k}:=\sum_{(s,a)}u^{(i)}_{k}(s,a)\left(T\rho^{\ast(i)}-\bar{r}^{(i)}(s,a)\right)$ that captures the difference between the $T$-step optimal reward $\rho^{\ast(i)}$ and the mean reward $\bar{r}^{(i)}(s,a)$ with $\sum_{k}u_{k}^{(i)}(s,a)=N_{k+1}^{(i)}(s,a)$ and $\sum_{(s,a)}N_{k}^{(i)}(s,a)=T$.
@@
where $\Delta_{k,M^{(i)}\in\mathcal{M}^{(i)}_{k}}^{r(i)}:=\sum_{s,a}u_{k}^{(i)}(s,a)\bigg{(}\tilde{r}^{(i)}_{k}(s,a)-\bar{r}^{(i)}(s,a)\bigg{)}$ and $\Delta_{k,M^{(i)}\in\mathcal{M}^{(i)}_{k}}^{p(i)}:=\boldsymbol{u}^{\top(i)}_{k}\bigg{(}\left(\tilde{\mathcal{P}}^{(i)}_{k}\left(\cdot\mid s,\tilde{\pi}_{k}(s)\right)\right)_{s}-\mathcal{I}\bigg{)}\boldsymbol{w}^{(i)}_{k}$. Besides, $\boldsymbol{u}^{(i)}_{k}:=\left(u_{k}^{(i)}(s,\tilde{\pi}_{k}(s))\right)_{s}$ denotes the vector representing the final count of state visits by agent $i$ in episode $k$, as determined under the policy $\tilde{\pi}_{k}(s)$, where the operator $(\cdot)_{s}$ concatenates the values for all $s\in\mathcal{S}$. The $\tilde{r}^{(i)}_{k}(s,a)$ are the maximal possible rewards according to a confidence interval similar to (11). $\left(\tilde{\mathcal{P}}^{(i)}_{k}\left(\cdot\mid s,\tilde{\pi}_{k}(s)\right)\right)_{s}$ is the transition matrix of the policy $\tilde{\pi}^{(i)}_{k}$ in the optimistic MDP $\tilde{M}^{(i)}_{k}$ computed via EVI and $\mathcal{I}$ typically represents the identity matrix. Furthermore, $\boldsymbol{w}_{k}^{(i)}$ denotes a modified value vector to indicate the state value range of EVI-iterated episode $k$ specific to agent $i$, where $w_{k}^{(i)}(s):=\eta_{k}^{(i)}(s)-\frac{1}{2}\left(\min_{s\in\mathcal{S}^{(i)}}\eta_{k}^{(i)}(s)+\max_{s\in\mathcal{S}^{(i)}}\eta_{k}^{(i)}(s)\right)$ and $\eta_{k}^{(i)}(s)$ are the state values given by EVI..
@@
where $C_{1}$ denotes a constant. Moreover, the term $\Delta_{k}^{p_{2}(i)}$ in (18) can be bounded by applying the Azuma-Hoeffding inequality
@@ V-B2 Analysis for Group Regret with Limited-Range Communications
Before analyzing the group regret in the multi-agent system within a constrained communication range, we start with some necessary notations and assume all CAVs constitute a graph. Considering the power graph $\mathcal{G}_{d,k}$, which is derived by selecting nodes from the original graph and connecting all node pairs in the new graph that are at a distance less than or equal to a specified value $d$ in the original graph. The neighborhood graph $\mathcal{G}_{d,k}^{(i)}=\left(\mathcal{F}_{d,k}^{(i)},\mathcal{E}^{(i)}_{d,k}\right)$ represents a sub-graph of $\mathcal{G}_{d,k}$, where $\mathcal{E}^{(i)}_{d,k}$ comprises the set of communication links connecting the agents $i$ and $i^{\prime}\in\mathcal{F}_{d,k}^{(i)}$. We also define the total number of state-action observations for agent $i$, which is calculated as the sum of observations from its neighbors within the communication range $d\in[0,D(\mathcal{G}_{k})]$ before episode $k$, denoted as $N_{d,k}^{(i)+}(s,a):=N_{d,k-1}^{(i)+}(s,a)+\sum_{i^{\prime}\in\mathcal{G}_{d,k}^{(i)}}u_{k}^{(i^{\prime})+}(s,a)$. Given the complication to directly estimate $N_{d,k}^{(i)+}(s,a)$, we introduce the concept of clique cover[24] for $\mathcal{F}_{d,k}^{(i)}$, which is a collection of cliques that can cover all vertices of a power graph. Further, let $\mathbf{C}_{d,k}$ denote a clique cover of $\mathcal{G}_{d,k}$ and a clique $\mathcal{C}\in\mathbf{C}_{d,k}$. Meanwhile, $|\mathcal{C}|$ represents the size of the clique $\mathcal{C}$. Additionally, the clique covering number $\bar{\chi}\left(\mathcal{G}_{d,k}\right)$ signifies the minimum clique number means finding a minimum number of cliques to cover the power graph $\mathcal{G}_{d,k}$ within the communication. Furthermore, characterized by a clique covering number $\bar{\chi}\left(\mathcal{G}_{d}\right):=\max_{k\in[K]}\{\bar{\chi}\left(\mathcal{G}_{d,k}\right)\}$, $\mathbf{C}_{d}$ constitutes the clique cover for the graph $\mathcal{G}_{d}$.
@@
In this setup, all agents within a clique $\mathcal{C}$ explore in proportion to the clique’s size and share the collected samples among themselves. For simplicity, we assume that $\mathcal{G}_{d,k}$ is connected, meaning that all cliques can communicate with one another. For simplicity, we take $N_{\mathcal{C},k}^{+}(s,a)$ to be the number of samples exchanged within the clique $\mathcal{C}$ before an episode $k$ that are available for all agents $i\in\mathcal{C}$. In the most challenging scenario, characterized by uniformly random exploration, the algorithm fails to capitalize on any inherent structures or patterns within the environment to potentially refine its learning approach. Consequently, this necessitates a maximum quantity of samples to ascertain an optimal policy, thus giving the worst case. Under these conditions, we present the following lemma.
@@
Based on the results of the single-agent regret bound described in Section V-B1, it is ready to analyse the upper bound on the group regret of the multi-agent setting under limited communication range, as shown in Theorem 1.
@@
Theorem 1 unveils the group regret bound and demonstrates that with respect to $T$, a sub-linear increase in the group regret bound is attained in (21). On the contrary in the standard non-communicative reinforcement learning setting for the same type of algorithm, the group regret is essentially equal to the sum of the single-agent regrets, as implied in Lemma 2.

### v2

@@ III-B System Model
$s.t.\quad\tilde{\mathcal{P}}^{(i)}\propto\mathcal{D}_{0:t}^{(i)+},0\leq\textit{d}\leq D(\mathcal{G}_{k})$ (7)
@@ IV-B Planning and Control with Learned Dynamics
Based on a learned dynamics model $\tilde{\mathcal{P}}^{(i)}$, MA-PETS tries to obtain a solution to (6) by resorting to a sample-efficiency controller MPC [25]. Generally speaking, MPC excels in providing high-precision trajectory planning, thus reducing the risk of collisions during vehicle operation. Moreover, it can predict both aleatoric and epistemic uncertainty accurately within the learned dynamics model [32]. Without loss of generality, assume that for a time-step $t$, agent $i$ observes the state $s_{t}^{(i)}$. Afterward, agent $i$ leverages the Cross-Entropy Method (CEM) [51] to generate $Q$ candidate action sequences within a $w$-horizon of MPC. Initially, each candidate action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)},\forall q\in[Q]$ is sampled following a Gaussian distribution $\mathcal{N}(\mu_{\phi},\sigma_{\phi})$ parameterized by $\phi$. Meanwhile, agent $i$ complements possible next-time-step states from the dynamics model ensembles, by adopting particle-based TS [52]. In other words, for each candidate action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)}$, TS predicts plausible state-action trajectories by simultaneously creating $P$ particles that propagate a set of Monte Carlo samples from the state $s_{t}^{(i)}$. Due to the randomness in the learned and assumed time-invariant dynamics model, each particle $p\in[P]$ can be propagated by $s_{t^{\prime}+1,p}^{(i)}={\tilde{\mathcal{P}}_{\theta_{b}}^{(i)}}\left(s_{t^{\prime},p}^{(i)},wherea_{t^{\prime}}^{(i)}\right)$ ($t^{\prime}\in[t,t+w-2]$), and $\tilde{\mathcal{P}}_{\theta_{b}}^{(i)},b\in[B]$ is a randomly selected dynamics model in the ensemble model. Therefore, the set of plausible state-action trajectories for each candidate action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)}=\left(a_{t}^{(q,i)},\ldots,a_{t+w-1}^{(q,i)}\right)$ ($q\in[Q]$) consists of $P$ parallel propagated states with the same action sequence $\mathbf{A}_{t:t+w-1}^{(q,i)}$. Afterward, given the calibrated known reward function $\mathcal{R}^{(i)}$, the evaluation of a candidate action sequence can be derived from the average cumulative reward of $P$ different parallel action-state trajectories, that is,
@@ V-A EVI for an Optimistic MDP
Notably, Assumption 1 holds naturally, if we neglect the possible discretization error of DNN in MA-PETS, as it has a trivial impact on understanding the sample efficiency incurred by information exchange. The quantization discussed here involves discretizing continuous state and action spaces. Common discretization methods [53] include uniform quantization, which divides each dimension of the continuous state space into intervals of equal width (bins). These intervals define the precision of discretization, with each interval representing a discrete state. Besides, non-uniform quantization may discretize based on the distribution or importance of states, and clustering algorithms (such as k-means or DBSCAN [54],[55]) can group points in the state space, with each cluster center representing a discrete state. Meanwhile, text encoding (e.g., Tile Coding [56]), which involves overlaying different tiles on the states and assigning one bit to each tile, belongs to an alternative solution. Therefore, if an agent is in a certain position, the corresponding discrete state can be represented by one bit-vector, with activated positions set to $1$ and all other positions set to 00. On the other hand, Assumption 2 could be easily met by intentionally ignoring the experienced visits of some “diligent” agents. Moreover, Assumption 3 implies unanimously scaling the group regret bound by $r_{\max}$, which does not affect learning the contributing impact from inter-agent communications.
@@
Consistent with UCRL2, we allow each agent $i$ to enter into a new episode $k+1$ once there exists a state-action pair $(s,a)$ that has just been played and satisfies $u_{k}^{(i)}(s,a)=N^{(i)+}_{k}(s,a)$. Here $u_{k}^{(i)}(s,a)$ denotes the number of state-action $(s,a)$ visits by agent $i$ in the episode $k$ and $N^{(i)+}_{k}(s,a)=\max\left\{1,N^{(i)}_{k}(s,a)\right\}$. By Assumption 2, all agents could enter into the next episode in a “simultaneous” manner, by intentionally ignoring the experienced visits of a “diligent” agent $i$ after meeting $u_{k}^{(i)}(s,a)=N^{(i)+}_{k}(s,a)$ in episode $k$. Therefore, at the very beginning of episode $k+1$, for each agent $i$, $N_{k+1}^{(i)}(s,a)=N_{k}^{(i)}(s,a)+u_{k}^{(i)}(s,a)$ for all state-action pairs$(s,a)$ This setting is similar to the doubling criterion in single-agent UCRL2 and ensures that each episode is long enough to allow sufficient learning. Furthermore, different from the single-agent UCRL2, at the end of an episode $k$, consistent with in MA-PETS, each agent $i$ has access to all the set of neighboring CAVs $\mathcal{F}_{d,k}^{(i)}$, so as to obtain the latest dataset $\mathcal{D}_{t_{k}-H:t_{k}}^{(j)}$, $\forall j\in\mathcal{F}_{d,k}^{(i)}$ and constitute $\mathcal{D}_{0:t_{k}}^{(i)+}$. Again, according to Assumption 2, $H=\sum_{s,a}u_{k}^{(i)}(s,a)$. Afterward, the EVI proceeds to estimate the transition probability function $\mathcal{P}^{(i)}(\cdot\mid s,a)$. In particular, various high-probability confidence intervals of the true MDP $M^{(i)}$ can be conjectured according to different concentration inequalities. For example, based on Hoeffding’s inequality [57], the confidence interval for the estimated transition probabilities can be given as
@@
Correspondingly, a policy for an optimistic MDP can be attained from learning in the conjectured plausible MDPs, and then the impact of the communication range $d$ on the performance of MA-PETS can be analyzed, in terms of the multi-agent group regret in (5).
@@ V-B1 Results for Single-Agent Regret Bound
Consider $M^{(i)}$ as a time-homogeneous and communicating MDP with a policy $\pi^{\prime\ast(i)}$. Meanwhile, by Lemma 1, the optimal $T$-step reward $\mathcal{R}^{\ast(i)}_{T}(M^{(i)})$ can be approximated by the optimistic average reward $\rho^{\ast(i)}$. Therefore, if $\pi^{\prime\ast}(i)$ is implemented in $M^{(i)}$ for $T$ steps, and we define the regret in a single episode $k$ as $\Delta^{(i)}_{k}:=\sum_{(s,a)}u^{(i)}_{k}(s,a)\left(T\rho^{\ast(i)}-\bar{r}^{(i)}(s,a)\right)$, which captures the difference between the $T$-step optimal reward $\rho^{\ast(i)}$ and the mean reward $\bar{r}^{(i)}(s,a)$, with $\sum_{k}u_{k}^{(i)}(s,a)=N_{k+1}^{(i)}(s,a)$ and $\sum_{(s,a)}N_{k}^{(i)}(s,a)=T$, we have the following lemma.
@@
Under Lemma 1, the differences between the observed rewards $r_{t}^{(i)}$, acquired when agent $i$ follows the policy $\pi^{\prime\ast(i)}$ to choose action $a_{t}^{(i)}$ in state $s_{t}^{(i)}$ at step $t$, and the aforementioned optimistic average reward $\rho^{\ast(i)}$ forms a martingale difference sequence. Besides, based on Hoeffding’s inequality, with a probability at least $1-\frac{\delta}{12(T)^{5/4}}$, the single-agent bound can be bounded as
@@
where $C_{1}=\frac{5}{8}$ is a constant.
@@
Define $\Delta_{k,M^{(i)}\in\mathcal{M}^{(i)}_{k}}^{r(i)}:=\sum_{s,a}u_{k}^{(i)}(s,a)\bigg{(}\tilde{r}^{(i)}_{k}(s,a)-\bar{r}^{(i)}(s,a)\bigg{)}$ as the accumulated regret in reward estimation over all state-action pairs. $\tilde{r}^{(i)}_{k}(s,a)$ denotes the maximal possible reward according to a confidence interval similar to (11). Meanwhile, $\Delta_{k,M^{(i)}\in\mathcal{M}^{(i)}_{k}}^{p(i)}:=\left(\boldsymbol{u}^{(i)}_{k}\right)^{\top}\bigg{(}\left(\tilde{\mathcal{P}}^{(i)}_{k}\left(\cdot\mid s,\tilde{\pi}_{k}(s)\right)\right)_{s}-\mathcal{I}\bigg{)}\boldsymbol{w}^{(i)}_{k}$ quantifies the regret from the estimation of transition probabilities. Additionally, $\boldsymbol{u}^{(i)}_{k}:=\left(u_{k}^{(i)}(s,\tilde{\pi}_{k}(s))\right)_{s}$ denotes the vector representing the final count of state visits by agent $i$ in episode $k$, as determined under the policy $\tilde{\pi}_{k}(s)$, where the operator $(\cdot)_{s}$ concatenates the values for all $s\in\mathcal{S}$. $\left(\tilde{\mathcal{P}}^{(i)}_{k}\left(\cdot\mid s,\tilde{\pi}_{k}(s)\right)\right)_{s}$ is the transition matrix of the policy $\tilde{\pi}^{(i)}_{k}$ in the optimistic MDP $\tilde{M}^{(i)}_{k}$ computed via EVI and $\mathcal{I}$ typically represents the identity matrix. Furthermore, $\boldsymbol{w}_{k}^{(i)}$ denotes a modified value vector to indicate the state value range of EVI-iterated episode $k$ specific to agent $i$, where $w_{k}^{(i)}(s):=\eta_{k}^{(i)}(s)-\frac{1}{2}\left(\min_{s\in\mathcal{S}^{(i)}}\eta_{k}^{(i)}(s)+\max_{s\in\mathcal{S}^{(i)}}\eta_{k}^{(i)}(s)\right)$ and $\eta_{k}^{(i)}(s)$ is the state value given by EVI. We can then decompose the aggregate regret considering $M^{(i)}\in\mathcal{M}^{(i)}_{k}$, which leads to the result in Lemma 4.
@@
where $C_{1}$ denotes a constant. Moreover, by applying the Azuma-Hoeffding inequality, the term $\Delta_{k}^{p_{2}(i)}$ in (18) can be bounded as
@@ V-B2 Analysis for Group Regret with Limited-Range Communications
Before analyzing the group regret in the multi-agent system within a constrained communication range, we start with some necessary notations and assume all CAVs constitute a graph. Considering the power graph $\mathcal{G}_{d,k}$, which is derived by selecting nodes from the original graph and connecting all node pairs in the new graph that are at a distance less than or equal to a specified value $d$ in the original graph. The neighborhood graph $\mathcal{G}_{d,k}^{(i)}=\left(\mathcal{F}_{d,k}^{(i)},\mathcal{E}^{(i)}_{d,k}\right)$ represents a sub-graph of $\mathcal{G}_{d,k}$, where $\mathcal{E}^{(i)}_{d,k}$ comprises the set of communication links connecting the agents $i$ and $i^{\prime}\in\mathcal{F}_{d,k}^{(i)}$. We also define the total number of state-action observations for agent $i$, which is calculated as the sum of observations from its neighbors within the communication range $d\in[0,D(\mathcal{G}_{k})]$ before episode $k$, denoted as $N_{d,k}^{(i)+}(s,a):=N_{d,k-1}^{(i)+}(s,a)+\sum_{i^{\prime}\in\mathcal{G}_{d,k}^{(i)}}u_{k}^{(i^{\prime})+}(s,a)$. Given the complication to directly estimate $N_{d,k}^{(i)+}(s,a)$, we introduce the concept of clique cover[27] for $\mathcal{F}_{d,k}^{(i)}$, which is a collection of cliques that can cover all vertices of a power graph. Further, let $\mathbf{C}_{d,k}$ denote a clique cover of $\mathcal{G}_{d,k}$ and a clique $\mathcal{C}\in\mathbf{C}_{d,k}$. Meanwhile, $|\mathcal{C}|$ represents the size of the clique $\mathcal{C}$. Additionally, the clique covering number $\bar{\chi}\left(\mathcal{G}_{d,k}\right)$ signifies the minimum number of cliques to cover the power graph $\mathcal{G}_{d,k}$ within the communication. Furthermore, characterized by a clique covering number $\bar{\chi}\left(\mathcal{G}_{d}\right):=\max_{k\in[K]}\{\bar{\chi}\left(\mathcal{G}_{d,k}\right)\}$, $\mathbf{C}_{d}$ constitutes the clique cover for the graph $\mathcal{G}_{d}$.
@@
In this setup, all agents within a clique $\mathcal{C}$ explore in proportion to the clique’s size and share the collected samples among themselves. For simplicity, we assume that $\mathcal{G}_{d,k}$ is connected, meaning that all cliques can communicate with one another. Meanwhile, we take $N_{\mathcal{C},k}^{+}(s,a)$ to be the number of samples exchanged within the clique $\mathcal{C}$ before an episode $k$ that are available for all agents $i\in\mathcal{C}$. In the most challenging scenario, characterized by uniformly random exploration, the algorithm fails to capitalize on any inherent structures or patterns within the environment to potentially refine its learning approach. Consequently, this necessitates a maximum quantity of samples to ascertain an optimal policy, thus giving the worst case. Under these conditions, we present the following lemma.
@@
Based on the results of the single-agent regret bound described in Section V-B1, it is ready to analyze the upper bound on the group regret of the multi-agent setting under limited communication range, as shown in Theorem 1.
@@
Theorem 1 unveils the group regret bound and demonstrates that concerning $T$, a sub-linear increase in the group regret bound is attained in (21). On the contrary, in the standard non-communicative reinforcement learning setting for the same type of algorithm, the group regret is essentially equal to the sum of the single-agent regrets, as implied in Lemma 2.